Solve: 7x-3y=26
2x+5y=25
: Solve: 7x-3y=26
2x+5y=25

Answer by jojo14344(1023)

Line Eqn's: 7x-3y=26 -----------> A
2x+5y=25 -----------------------> B
In A:
3y=7x-26
cross(3)y/cross(3)=(7x-26)/3
y=(7/3)x-(26/3) ----------------> C
Subst. C in B:
2x+5((7/3)x-(26/3))=25
2x+(35/3)x-(130/3)=25
(6x+35x)/3=25+(130/3)=(75+130)/3=205/3,
(41x)/cross(3)=205/cross(3)
cross(41)x/cross(41)=cross(205)5/cross(41)
highlight(x=5)
Subst. in C:
y=(7/3)5-(26/3)=(35/3)-(26/3)=9/3
highlight(y=3)
See the graph:
drawing(400,400,-6,15,-10,10,grid(1),graph(400,400,-6,15,-10,10,(-2/5)x+5,(7/3)x-(26/3)),circle(5,3,.20))) -------> Line A (Green) intersect at Line B (RED) at points (5,3) as shown (as computed too)
.
Thank you,
Jojo